OpenStudy (jkbo):

Please help me 3m/5 + a = b − 2m/5 A.5/3a − b B.3/5a − b C.3/5b − a D.m = b − a

OpenStudy (anonymous):

add 2m/5 to both sides

OpenStudy (anonymous):

\[\frac{ 3m }{ 5 } + a = b - \frac{ 2m }{ 5 }\] Add \[\frac{ 2m }{ 5 }\] to both sides \[\frac{ 3m }{ 5 } + a +\frac{ 2m }{ 5 }\ = b - \frac{ 2m }{ 5 }\ +\frac{ 2m }{ 5 }\] \[-\frac{ 2m }{ 5 }\ +\frac{ 2m }{ 5 }\] cancel out Thus \[\frac{ 5m }{ 5 } + a = b \] Divide 5m by 5 \[m + a = b \]

OpenStudy (anonymous):

yes :)

OpenStudy (jkbo):

thank you both

OpenStudy (jkbo):

I got 1 @edyscape

OpenStudy (jkbo):

@mathessentials I got 1

OpenStudy (anonymous):

explain

OpenStudy (jkbo):

5m/5 I got just 1

OpenStudy (anonymous):

5m/5 = m

OpenStudy (anonymous):

you can cancel the 5 you CANT cancel the m - it's only in the numerator!

OpenStudy (anonymous):

\[\frac{ 5m }{ 5 } =\frac{ 5 * m }{ 5 } = 1 * m = 1m = m \] If you want 1 then you have to do this \[\frac{ 5m }{ 5m } =\frac{ 5 * m }{ 5 * m } = 1 * 1 = 1 \]

OpenStudy (anonymous):

the answer is d

OpenStudy (jkbo):

@edyscape so it would be D.? @killerdime20

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yes, it is D because \[m + a = b\] \[m + a -a = b - a\] \[m = b - a\]